Write the equation for the agent’s budget constraint between consumption and leisure when he has positive wealth.
Answer #1

Assume that there is a wealth Z available with an agent. Fraction of the working hours is given by 1 - L where L is the fraction of leisure hours. Considering the wage rate 'w', consumption is given by C = (1 - L)*w + Z.

Rearrange this equation to get the budget constraint

C + wL = w + Z